### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) models are critically undermined by citation hallucinations, a deceptive failure where a model confidently cites a source that fails to support its claim. Existing work often attributes hallucination to a simple over-reliance on the model’s parametric knowledge. We challenge this view and introduce FACTUM (Framework for Attesting Citation Trustworthiness via Underlying Mechanisms), a framework of four mechanistic scores measuring the distinct contributions of a model’s attention and FFN pathways, and the alignment between them. Our analysis reveals two consistent signatures of correct citation: a significantly stronger contribution from the model’s parametric knowledge and greater use of the attention sink for information synthesis. Crucially, we find the signature of a correct citation is not static but evolves with model scale. For example, the signature of a correct citation for the Llama-3.2-3B model is marked by higher pathway alignment, whereas for the Llama-3.1-8B model, it is characterized by lower alignment, where pathways contribute more distinct, orthogonal information. By capturing this complex, evolving signature, FACTUM outperforms state-of-the-art baselines by up to 37.5% in AUC. Our findings reframe citation hallucination as a complex, scale-dependent interplay between internal mechanisms, paving the way for more nuanced and reliable RAG systems.
@@ 1 Introduction
Figure 1: An example demonstrating how citation hallucination makes factually correct information dangerously unreliable. Both scenarios receive the same query and retrieved documents. Left (Correct Citation): The model correctly cites its source ‘[Source: 1]’, making the claim verifiable. Right (Hallucinated Citation): The model generates the same correct fact but falsely attributes it to ‘[Source: 2]’, pointing the user to conflicting information and making the claim unverifiable. This isolates citation hallucination as a critical failure of the model’s attribution mechanism.
Retrieval-Augmented Generation (RAG) systems are designed to improve the accuracy of Large Language Models (LLMs) by integrating relevant information retrieved from external knowledge sources [21, 2]. However, even with access to accurate information, these models can still produce statements that misrepresent or contradict the provided sources, a phenomenon known as RAG Hallucination [51, 47]. Prior work has studied the use of external context and an LLM’s parametric knowledge to detect general RAG hallucinations [47, 48, 25]. In contrast, the critical issue of citation hallucination detection has been largely unaddressed. We argue that detecting hallucination of citations in RAG is a distinct problem that deserves its own study.
@@
This becomes particularly challenging in long-form RAG, where LLMs are tasked with generating extended, multi-paragraph reports from vast document collections [^1] Modern long-context models can process context windows from 32K to 2M tokens, equivalent to hundreds or thousands of pages of text [23].. In this scenario, the model must maintain correct attribution throughout a long output, introducing the risk of attributional drift: the longer the model writes, the more likely it is to forget which statement came from which source, leading to misattribution [6]. To identify the internal signatures that characterize hallucination of citations, our work moves beyond black-box analysis to investigate the measurable mechanistic signals within the model’s internal states.
@@
To locate these signals, we ground our approach in the transformer’s architecture. Each layer processes information through two distinct sub-modules with fundamentally different roles. The Attention mechanism serves as the sole pathway for retrieving information from the external context, processing the provided source documents [22, 7]. In contrast, the Feed-Forward Networks (FFNs) are the primary pathway for applying the model’s stored, parametric knowledge [13, 42]. This architectural division allows us to conceptualize them as two primary information streams: what the model reads and what it recalls. By decoupling their contributions to the model’s residual stream, which is the central information space where updates are written, we can directly probe the interplay between what the model reads from the external context and what it recalls from parametric memory [9, 29].
@@
Prior mechanistic analyses of general RAG hallucination argued that hallucinations predominantly arise from an over-reliance on the model’s parametric knowledge [47]. We argue that for the task of citation, this view is incomplete. The core failure is not a simple over-reliance on parametric knowledge, but a systemic breakdown in how the model utilizes its internal mechanisms. To reliably detect citation hallucination, we must ask a series of interconnected questions about the model’s state at the moment of citation generation. Regarding what the model reads via its Attention pathway: how strongly is the citation grounded in the source documents, and how is that information synthesized? Regarding what the model recalls from its FFNs: how forcefully does it apply its own parametric knowledge? Finally, how do these two pathways interact? Are they working in concert, or are they contributing distinct, orthogonal information?
@@
To answer these questions, our contributions are threefold: (1) we introduce FACTUM (Framework for Attesting Citation Trustworthiness via Underlying Mechanisms), a novel diagnostic framework using refined and novel mechanistic scores to measure these signals; (2) we demonstrate that by analyzing the complete signature of citation generation, FACTUM significantly outperforms state-of-the-art baselines for citation hallucination detection in long-form RAG; and (3) we provide empirical evidence challenging the conventional view, showing that strong FFN Pathway activity is constructive, not detrimental, to correct citation generation.
@@ 2 Related Work
Methods for detecting hallucinations in RAG systems largely fall into two categories: those that treat the model as a black box, and those that analyze its internal states. Black-box approaches evaluate the model’s final output without considering its internal processing [45]. This is often done by using a powerful external LLM as a judge [41, 10], fine-tuning a separate detector model [38, 44, 40, 19], or checking for consistency across multiple sampled outputs [31]. The primary limitation of these approaches is that they are purely behavioral [16, 46]. They can flag a potential hallucination, but they cannot reveal the underlying mechanistic signals within the model that give insight into the failure. In contrast, methods that analyze model internals seek measurable signals of hallucination [39]. Early approaches focused on general uncertainty, using signals such as token probabilities and entropy [52, 17, 30]. The most recent prior work is ReDeEP, which took a more mechanistic approach [47]. It was the first to decouple the model’s reliance on external context versus its internal parametric knowledge, proposing the External Context Score (ECS) and Parametric Knowledge Score (PKS). Our work adopts this decoupling framework but offers critical advancements: we introduce more direct and robust scores for each pathway, and add two novel dimensions ReDeEP does not address: a direct measure of pathway alignment and the role of the attention sink.
While foundational, ReDeEP’s scores have critical limitations for detecting citation hallucination. First, ECS is an imprecise measure of contextual grounding for this task. As it considers the entire prompt, it dilutes the crucial signal from source documents with noise from system instructions and the user query. This design also conflates two distinct attention signals: attention to the actual evidence and attention to the frequently utilized “attention sink” (beginning-of-sentence token), making it impossible to study their separate roles. Second, PKS relies on LogitLens to project hidden states into a vocabulary space, a proxy that has been shown to be unreliable [5]. Finally, ReDeEP offers no way to directly measure the geometric alignment or conflict between the model’s information pathways. FACTUM is designed to overcome these specific limitations.
@@
While hallucination in RAG is a broad topic, citation hallucination is a specific and critical sub-problem [8]. Recent work has established its prevalence, confirming that even state-of-the-art models frequently misattribute or fabricate sources [28]. Existing studies have approached this from a black-box perspective, for instance by asking an LLM follow-up questions to test its comprehension of a source [1, 37]. However, this still fails to answer the core mechanistic question: Is there a measurable signature within the model’s internal information pathways that reliably distinguishes a correct citation from a hallucinated citation? Our work is the first to systematically search for this internal signature, connecting citation hallucination to a predictable pattern in the model’s internal state.
Figure 2: An illustration of the FACTUM framework analyzing a model’s internal state at a citation token (e.g., 1), inspired by prior work [47]. The framework derives scores from the two primary pathways. From the Attention Pathway, it computes the Contextual Alignment Score (CAS) to measure grounding in the source documents, and the Beginning-of-Sentence Attention Score (BAS) to measure information synthesis. It also measures the FFN Pathway’s Parametric Force Score (PFS) to quantify the magnitude of the FFN’s update, and the Pathway Alignment Score (PAS) to assess the geometric alignment between the two pathway updates.
@@ 3 The FACTUM Framework
The FACTUM framework leverages the internal architecture of the Transformer to provide a direct, mechanistic view into the model’s state during citation generation. Rather than treating the model as a black box, we decompose its internal processing into two distinct information pathways: the Attention pathway (the reader) and the FFN pathway (the recall). By decoupling these updates to the residual stream, we define four mechanistic scores that are designed to overcome the limitations of prior methods by providing more direct, robust, and targeted measurements of the model’s internal processing. Together, these scores allow us to test our central hypothesis that citation hallucination arises from a systemic failure in the coordination between contextual grounding, internal synthesis, parametric force, and pathway alignment. Figure 2 provides a visual representation of these scores and how they are calculated.
FACTUM is grounded in the mechanistic interpretability view of Transformers, treating hidden states as points in a structured high-dimensional residual stream $\mathbf{x}\in\mathbb{R}^{d}$ [9]. Following prior work, we model the architecture as a shared communication channel where each sub-layer performs an additive write to that state. For a citation token at position $i$, the state at layer $l$ is the sum of the previous state and the updates from the multi-head attention (MHA) and feed-forward network (FFN) sub-layers:
@@
This linear, additive structure allows us to decouple the model’s internal processing into two primary information pathways, measuring their individual contributions to the communication channel [15, 29].
@@ 3.1 The Attention Pathway
The Attention Pathway serves as the model’s reader and is the sole mechanism for moving information between token positions [15, 50]. In RAG, its purpose is to contextualize the current token by selectively retrieving information from the source documents $T_{C}$ [22, 7]. Mechanistically, this pathway facilitates a targeted transport of features: it uses the QK circuit to determine the attention pattern $A_{i,j}$ (where to look) and the OV circuit ($W_{OV}$) to extract specific semantic content (what to read) [9]. The resulting update $\mathbf{v}_{\text{attn}}$ aggregates this retrieved evidence into the residual stream:
@@
To overcome the signal dilution of prior methods that consider the entire prompt, CAS provides a precise measure of contextual grounding by focusing exclusively on the source document tokens ($T_{C}$). It computes an attention-weighted average of the final-layer hidden states ($\mathbf{h}^{(L)}$) of all context tokens to form a holistic context vector, reflecting the model’s understanding of the evidence. This weighting by attention ensures that the most attended-to context tokens have the greatest influence. The score is the cosine similarity between this context vector and the generated citation token’s own final-layer hidden state:
@@
A high CAS indicates strong semantic alignment between the generated token and the attended-to evidence, serving as a direct measure of contextual grounding.
@@
A higher BAS suggests greater reliance on this internal synthesis mechanism, a signal entirely missed by previous approaches. For models without such a token, it measures attention to the very first token of the input, which serves the same functional role.
@@ 3.2 The FFN Pathway
The FFN Pathway represents the model’s recall mechanism and serves as the primary channel for injecting parametric knowledge into the residual stream [13, 7]. FFNs function as key-value memories where factual associations are stored and retrieved [11]. Following the key-value recall model, the FFN update vector at layer $l$ is defined as:
@@
where $W_{in}$ acts as keys that detect semantic patterns in the communication channel and $W_{out}$ acts as values that inject the associated parametric information [11]. Crucially, direct interventions on FFNs have been shown to manipulate the factual content of a model’s output [34, 35, 24]. By measuring the activity of these memories, we can quantify the intensity of the model’s internal updates during citation generation.
@@
A larger norm signifies a more substantial modification to the residual stream, indicating a stronger “force” from the model’s stored knowledge.
@@ 3.3 Pathway Interaction
A key component of our hypothesis is that citation hallucination arises from a systemic coordination failure across all four of our measured signals is the direct geometric alignment between the model’s information pathways. While not the sole indicator, this alignment provides a crucial signal about whether the model’s parametric knowledge is working with or against the contextual evidence. The Pathway Alignment Score (PAS) is the first metric designed to explicitly quantify this specific interaction.
@@
The PAS metric quantifies the geometric alignment between the attention and FFN pathways by computing the direct cosine similarity between their respective update vectors. This provides a rich signal that distinguishes between collaboration, conflict, and uncoordinated activity. The score is calculated as:
@@
The attention update vector, $\mathbf{v}_{\text{attn}}^{(l)}$, is the change to the residual stream from the attention mechanism, where $\mathbf{x}_{\text{input}}^{(l)}$ and $\mathbf{x}_{\text{pre-ffn}}^{(l)}$ are the states of the residual stream before and after the attention block, respectively. The score’s value lies in its full range: scores near +1+1 indicate alignment (collaboration), scores near 0 indicate orthogonality (unrelated processing), and scores near 1-1 indicate anti-alignment where the FFN actively negates the contextual information.
@@ 4.2 Dataset
Our work focuses on detecting citation hallucination in a challenging long-form, multi-document RAG scenario. To this end, we use the TREC NeuCLIR 2024 dataset [^2] Dataset available at: https://neuclir.github.io/ [20]. Its report generation task, requiring synthesizing information from retrieved documents, is significantly more challenging than the single-passage QA tasks often used in prior RAG hallucination studies. In our experiments on NeuCLIR, we use 15 documents as input to the generator. This multi-document scenario creates a high risk of the “attributional drift” that our work aims to detect. Although NeuCLIR is a cross-lingual track, our analysis is performed in a monolingual English setting: we provide the models with machine-translated English versions of the source documents to generate English reports.
Crucially, while our experimental setup uses this 15-document implementation, the FACTUM framework is agnostic to the retrieval system. It operates at the moment of generation, making its scores applicable to any RAG system, regardless of the number or quality of retrieved documents.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) models are critically undermined by citation hallucinations, a deceptive failure where a model cites a source that fails to support its claim. While existing work attributes hallucination to a simple over-reliance on parametric knowledge, we reframe this failure as an evolving, scale-dependent coordination failure between the Attention (reading) and Feed-Forward Network (recalling) pathways. We introduce FACTUM (Framework for Attesting Citation Trustworthiness via Underlying Mechanisms), a framework of four mechanistic scores: Contextual Alignment (CAS), Attention Sink Usage (BAS), Parametric Force (PFS), and Pathway Alignment (PAS). Our analysis reveals that correct citations are consistently marked by higher parametric force (PFS) and greater use of the attention sink (BAS) for information synthesis. Crucially, we find that “one-size-fits-all” theories are insufficient as the signature of correctness evolves with scale: while the 3B model relies on high pathway alignment (PAS), our best-performing 8B detector identifies a shift toward a specialized strategy where pathways provide distinct, orthogonal information. By capturing this complex interplay, FACTUM outperforms state-of-the-art baselines by up to 37.5% in AUC. Our results demonstrate that high parametric force is constructive when successfully coordinated with the Attention pathway, paving the way for more nuanced and reliable RAG systems.
@@ 1 Introduction
Figure 1: An example demonstrating how citation hallucination makes factually correct information dangerously unreliable. Both scenarios receive the same query and retrieved documents. Left (Correct Citation): The model correctly cites its source ‘[Source: 1]’, making the claim verifiable. Right (Hallucinated Citation): The model generates the same correct fact but falsely attributes it to ‘[Source: 2]’, pointing the user to conflicting information and making the claim unverifiable.
Retrieval-Augmented Generation (RAG) systems are designed to improve the accuracy of Large Language Models (LLMs) by integrating relevant information retrieved from external knowledge sources [21, 2]. However, even with access to accurate information, these models can still produce statements that misrepresent or contradict the provided sources, a phenomenon known as RAG Hallucination [51, 46]. While prior work leverages external context and parametric knowledge to detect general RAG hallucinations [46, 47, 25], the critical issue of citation hallucination detection has been largely unaddressed. We argue that detecting citation hallucinations in RAG is a distinct problem that deserves its own study.
@@
This becomes particularly challenging in long-form RAG, where LLMs are tasked with generating extended, multi-paragraph reports from vast document collections [^1] Modern long-context models can process context windows from 32k to 2M tokens, equivalent to hundreds or thousands of pages of text [23].. In this setting, the model must maintain correct attribution throughout a long output, introducing the risk of attributional drift: the longer the model writes, the more likely contextual signals regarding source attribution dissipate, leading to misattribution [6]. To identify the internal signatures that characterize hallucination of citations, our work moves beyond black-box analysis to investigate the measurable mechanistic signals within the model’s internal states.
@@
To locate these signals, we ground our approach in the transformer’s architecture, where each layer processes information through two distinct submodules with fundamentally different roles. The attention mechanism serves as the pathway to retrieve information from the external context, processing the source documents provided [22, 7]. In contrast, Feed-Forward networks (FFNs) are the primary pathway to apply the stored parametric knowledge of the model [14, 41]. This architectural division allows us to conceptualize them as two primary information pathways: what the model reads and what it recalls. By decoupling their contributions to the model’s residual stream, which is the shared communication channel that sub-layers both read from and write updates to, we can directly probe the interplay between what the model reads from the external context and what it recalls from parametric memory [9, 29].
@@
Prior mechanistic analyses of general RAG hallucination argued that hallucinations predominantly arise from an over-reliance on the model’s parametric knowledge [46]. We argue that for citation attribution, this view is incomplete. The core failure is not a simple over-reliance on parametric knowledge, but a systemic breakdown in how the model utilizes its internal mechanisms. Detecting such failures reliably necessitates evaluating the model’s mechanistic state during citation generation. To this end, we ask a series of interconnected questions: Regarding what the model reads via its Attention pathway: How strongly is the citation grounded in the source documents, and how is that information synthesized across the model’s layers? Regarding what the model recalls from its FFNs: How forcefully does it apply its own parametric knowledge? Finally, how do these two pathways interact? Are they working in concert, or are they contributing distinct, orthogonal information?
@@
To answer these questions, our contributions are threefold: (1) we introduce FACTUM (Framework for Attesting Citation Trustworthiness via Underlying Mechanisms), a novel diagnostic framework designed to capture scale-dependent coordination between internal pathways utilizing four mechanistic scores to measure these signals; (2) we demonstrate that by analyzing the citation signature, FACTUM significantly outperforms state-of-the-art baselines for citation hallucination detection in long-form RAG; and (3) we provide empirical evidence challenging the conventional view, showing that strong FFN Pathway activity is constructive, not detrimental, to correct citation generation.
@@ 2 Related Work
Methods for detecting hallucinations in RAG largely fall into two categories: those that treat the model as a black box, and those that analyze its internal states. Black-box approaches evaluate the model’s final output without considering its internal processing [44]. Techniques include using an external LLM judge [40, 10], fine-tuning a separate detector model [37, 43, 39, kovács2025lettucedetecthallucinationdetectionframework], or checking for consistency across sampled outputs [31]. The primary limitation of these approaches is that they are purely behavioral [17, 45]. They can flag a potential hallucination, but they cannot reveal the underlying mechanistic signals within the model that give insight into the failure. In contrast, methods that analyze model internals seek measurable signals of hallucination [38]. While early approaches focused on general uncertainty, leveraging signals such as token probabilities and entropy [52, 18, 30], the most relevant prior work is ReDeEP which introduced a framework for decoupling a model’s reliance on external context versus parametric knowledge [46]. However, ReDeEP treats these pathways as independent, monolithic entities and relies on imprecise proxies that significantly limit its effectiveness for detecting citation hallucination.
Specifically, ReDeEP’s External Context Score (ECS) considers the entire prompt, diluting the context signal from source documents with noise from system instructions and user queries. This design also conflates attention to actual evidence with attention to the functional “attention sink” (beginning-of-sentence token), preventing role isolation. Furthermore, its Parametric Knowledge Score (PKS) relies on LogitLens, a vocabulary-space projection that is an unreliable proxy for the residual stream’s true state [5]. FACTUM overcomes these limitations by introducing more direct, targeted scores for each pathway. Beyond mere decoupling, we move toward a coordination-based framework that introduces the role of the attention sink (BAS) and the first direct measure of geometric alignment between pathways (PAS). This allows FACTUM to identify the scale-dependent coordination failures that monolithic baselines miss.
@@
While hallucination in RAG is a broad topic, citation hallucination is a specific and critical sub-problem [8]. Recent work has established its prevalence, confirming that even state-of-the-art models frequently misattribute or fabricate sources [28]. Existing studies have approached this from a black-box perspective, for instance by asking an LLM follow-up questions to test source comprehension [1, 36]. However, this fails to address the core mechanistic question: Is there a measurable signature within the model’s internal information pathways that reliably distinguishes a correct citation from a hallucinated citation? Our work is the first to characterize this internal signature, linking citation hallucination to predictable patterns in the model’s internal states.
@@ 3 The FACTUM Framework
Figure 2: The FACTUM framework analyzing internal states at a citation token (e.g., 1), building on [46]. It derives scores from two pathways: from the Attention Pathway, it computes the Contextual Alignment Score (CAS) to measure grounding in the source documents, and the Beginning-of-Sentence Attention Score (BAS) to measure information synthesis; from the FFN pathway it measures Parametric Force Score (PFS) to quantify the magnitude of the FFN’s update, and the Pathway Alignment Score (PAS) to assess the geometric alignment between the two pathway updates.
The FACTUM framework leverages the Transformer architecture to provide a direct, mechanistic view into the model’s internal states during citation generation. This approach is grounded in the mechanistic interpretability view of Transformers, treating hidden states as points in a structured, high-dimensional residual stream $\mathbf{x}\in\mathbb{R}^{d}$ [9]. Following prior work, we treat this stream as a shared communication channel where each sub-layer reads from the preceding state and performs an additive write to update it [16, 29]. For a citation token at position $i$, the state at layer $l$ is defined as the sum of the previous state and the updates from the Multi-Head Attention (MHA) and Feed-Forward Network (FFN) sub-layers:
@@
This linear, additive structure allows us to move beyond black-box analysis by decomposing internal processing into two primary information pathways: the Attention pathway (read) and the FFN pathway (recall). By measuring their individual contributions to the communication channel, we define four mechanistic scores designed to provide more direct, robust, and targeted measurements than prior methods. Together, these scores allow us to test our central hypothesis that citation hallucination arises from a systemic failure in the coordination between contextual grounding, internal synthesis, parametric force, and pathway alignment. Figure 2 provides a visual representation of these scores and how they are calculated.
@@ 3.1 The Attention Pathway
The Attention Pathway acts as the model’s reader and is the sole mechanism for moving information between token positions [16, 49]. In RAG, its purpose is to contextualize the current token by selectively retrieving information from the source documents $T_{C}$ [22, 7]. Mechanistically, this pathway performs targeted feature transport: it uses the QK circuit to determine the attention pattern $A_{i,j}$ (where to look) and the OV circuit ($W_{OV}$) to extract specific semantic content (what to read) [9]. The resulting update $\mathbf{v}_{\text{attn}}$ aggregates this retrieved evidence into the residual stream:
@@
To overcome the signal dilution of prior methods that consider the entire prompt, CAS provides a precise measure of contextual grounding by focusing exclusively on the source document tokens ($T_{C}$). It computes an attention-weighted average of the final-layer hidden states ($\mathbf{h}^{(L)}$) of all context tokens to form a holistic context vector, reflecting the model’s representation of the evidence. This weighting by attention ensures that the most attended-to context tokens have the greatest influence. The score is the cosine similarity between this context vector and the generated citation token’s own final-layer hidden state:
@@
where $\mathbf{h}_{t_{j}}^{(L)}$ represents the final-layer hidden state of context token $j$, and $A_{i,j}^{(h,l)}$ is the attention weight from head $h$ in layer $l$. A high CAS indicates strong semantic alignment between the generated token and the attended-to evidence, serving as a direct measure of contextual grounding.
@@
where $A_{i,1}^{(l,h)}$ is the attention weight from head $h$ in layer $l$ allocated to the first token (<s>) when generating citation $t_{i}$. A higher BAS suggests greater reliance on this internal synthesis mechanism, a signal entirely missed by previous approaches. For models without such a token, it measures attention to the very first token of the input, which serves the same functional role.
@@ 3.2 The FFN Pathway
The FFN Pathway acts as the recall mechanism, the primary channel for injecting parametric knowledge into the residual stream [14, 7]. FFNs function as key-value memories where factual associations are stored and retrieved [12]. Following the key-value model, the FFN update vector at layer $l$ is defined as:
@@
where $W_{in}$ acts as keys that detect semantic patterns in the communication channel and $W_{out}$ acts as values that inject the associated parametric information [12]. Crucially, direct interventions on FFNs have been shown to manipulate the factual content of a model’s output [33, 34, 24]. Measuring this activity allows us to quantify the parametric force behind each generated citation.
@@
Here, $||\cdot||_{2}$ denotes the L2 norm, and $\mathbf{x}_{\text{pre-ffn}}^{(l)},\mathbf{x}_{\text{post-ffn}}^{(l)}$ represent the states of the residual stream before and after the FFN block. A larger norm signifies a more substantial modification to the residual stream, indicating a stronger “force” from the model’s stored parametric knowledge.
@@ 3.3 Pathway Interaction
Central to our hypothesis is that citation hallucination stems from a systemic failure in coordination, measured via the geometric alignment between information pathways. While not the sole indicator, this alignment provides a rich signal of whether the model’s parametric recall reinforces or contradicts the contextual evidence. To capture this interplay, we introduce the Pathway Alignment Score (PAS), the first metric designed to explicitly quantify this interaction.
@@
The PAS metric quantifies the geometric alignment between the attention and FFN pathways by computing the cosine similarity between their respective update vectors. This provides a rich signal that distinguishes between collaboration, conflict, and uncoordinated activity. The score is calculated as:
@@
where $\mathbf{v}_{\text{attn}}^{(l)}$ is the attention update vector, which is the update from the attention block to the residual stream, calculated as the difference between the state after the attention block ($\mathbf{x}_{\text{pre-ffn}}$) and the state before it ($\mathbf{x}_{\text{input}}$). Scores near +1+1 indicate alignment (collaboration), scores near 0 indicate orthogonality (unrelated processing), and scores near 1-1 indicate anti-alignment where the FFN actively negates the contextual information.
@@ 4.2 Dataset
Our work focuses on detecting citation hallucination in a challenging long-form, multi-document RAG scenario. To this end, we use the TREC NeuCLIR 2024 dataset [^2] Dataset available at: https://neuclir.github.io/ [20]. Its report generation task, requiring synthesizing information from retrieved documents, is significantly more challenging than the single-passage QA tasks often used in prior RAG hallucination studies. In our experiments on NeuCLIR, we use 15 documents as input to the generator. This multi-document scenario creates a high risk of the “attributional drift” that our work aims to detect. Although NeuCLIR is a cross-lingual track, our analysis is performed in a monolingual English setting: we provide the models with machine-translated English versions of the source documents to generate English reports. Crucially, while our experimental setup uses this 15-document implementation, the FACTUM framework is agnostic to the retrieval system. It operates at the moment of generation, making its scores applicable to any RAG system, regardless of the number or quality of retrieved documents.